### v1

# Inv-Entropy: A Fully Probabilistic Framework for Uncertainty Quantification in Language Models

Abstract
Large language models (LLMs) have transformed natural language processing, but their reliable deployment requires effective uncertainty quantification (UQ). Existing UQ methods are often heuristic and lack a probabilistic foundation. This paper begins by providing a theoretical justification for the role of perturbations in UQ for LLMs. We then introduce a dual random walk perspective, modeling input–output pairs as two Markov chains with transition probabilities defined by semantic similarity. Building on this, we propose a fully probabilistic framework based on an inverse model, which quantifies uncertainty by evaluating the diversity of the input space conditioned on a given output through systematic perturbations. Within this framework, we define a new uncertainty measure, Inv-Entropy. A key strength of our framework is its flexibility: it supports various definitions of uncertainty measures, embeddings, perturbation strategies, and similarity metrics. We also propose GAAP, a perturbation algorithm based on genetic algorithms, which enhances the diversity of sampled inputs. In addition, we introduce a new evaluation metric, Temperature Sensitivity of Uncertainty (TSU), which directly assesses uncertainty without relying on correctness as a proxy. Extensive experiments demonstrate that Inv-Entropy outperforms existing semantic UQ methods. The code to reproduce the results can be found at https://github.com/UMDataScienceLab/Uncertainty-Quantification-for-LLMs.
1 Introduction
Large language models (LLMs) have demonstrated remarkable success in various natural language processing tasks, such as text generation, question answering, and summarization (Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023). These models, particularly driven by recent advances in transformers and attention mechanisms, have pushed the boundaries of what is achievable in language understanding and generation (Chung et al., 2024; OpenAI, 2023). However, despite their impressive capabilities, a significant challenge remains: LLMs tend to hallucinate, or generate confidently wrong predictions (Maynez et al., 2020; Zhang et al., 2024b). This is a serious concern in applications where reliability is paramount, such as healthcare, autonomous systems, and legal domains, where incorrect outputs can have dire consequences (Ji et al., 2023). Addressing these challenges is essential for ensuring the reliable and responsible deployment of LLMs at scale, ultimately unlocking their full potential in real-world applications.
A central step in addressing these limitations is developing effective measures for UQ, enabling LLMs to acknowledge their confidence in a generated output. Existing UQ approaches rely predominantly on heuristic consistency checks. Most commonly, they use the model’s own generation likelihood or perplexity as a proxy for confidence, or they measure dispersion across multiple sampled continuations (e.g. via n-gram overlap or embedding-space variance) (Mudumbai and Bell, 2024). Such likelihood and sampling-based measures, however, lack a grounded probabilistic interpretation. Also, token-level probabilities are known to dramatically under-estimate uncertainty as models can be “confidently wrong” (Jiang et al., 2021), and they are often inaccessible in black-box LLMs.
To probe deeper model brittleness, recent work has turned to input-perturbation UQ methods. Variant prompts are created through paraphrasing, adversarial token insertion, or temperature shifts. Output sensitivity is then quantified as an uncertainty signal (Gao et al., 2024; Tuna et al., 2022; Seeböck et al., 2019). For instance, Gao et al. (2024) randomly perturb both prompt wording and sampling temperature, then aggregate variation across outputs to flag unstable predictions. Tuna et al. (2022) applies adversarial paraphrases to uncover “blind spots” where small semantic-preserving edits trigger large output changes, while Seeböck et al. (2019) uses systematic character-level and word-level corruptions to map regions of high model vulnerability.
Despite these advances, a fundamental probabilistic gap remains, as existing methods mainly rely on heuristic score functions for UQ. To address this, we introduce Inv-Entropy, a fully probabilistic framework grounded in random-walk theory that offers a new perspective: it learns the distribution of possible inputs that could have produced a given outcome. This is accomplished through structured perturbations, which simultaneously capture input variability and the influence of different inputs on model predictions. Our contributions are summarized as follows:
  1. 1.
    We present the first work in UQ for LLMs that adopts a fully probabilistic framework, grounded in random walk theory. This framework is highly flexible, and its probabilistic nature allows for the use of various UQ measures. It is also intrinsically capable of handling black-box models, as it is agnostic to token probabilities.
  2. 2.
    We introduce an inverse perspective that quantifies input diversity given an output, inspired by “Asymmetry in Semantic Variability” (defined later). Extensive simulations highlight its advantages, especially for short-form answers where traditional methods often struggle.
  3. 3.
    Theoretically, our work provides a rigorous foundation for using perturbation-based methods for UQ, validating the recent momentum behind such approaches.
  4. 4.
    We propose GAAP, a novel perturbation algorithm that enhances input sampling diversity. Empirical results show that GAAP significantly improves perturbation-based UQ.
  5. 5.
    Finally, we introduce a new evaluation metric, Temperature Sensitivity of Uncertainty (TSU) capable of evaluating uncertainty without relying on correctness as a proxy. This enables evaluation of UQ on any dataset, even when labels are unavailable.
For the remainder of the paper, we use $\mathbb{E}[\cdot]$ to denote expectation and $\mathbb{V}[\cdot]$ to denote variance. We note that a more detailed related work section can be found in Appendix A.
2 Perturb-then-Quantify
2.1 Why perturb the input?
Unlike model-free UQ approaches (Shafer and Vovk, 2008), we exploit a key property specific to language: semantic equivalence. An ideal LLM should respond similarly to semantically equivalent inputs. Therefore, we can design an approach based on input perturbations on the semantic equivalent class to detect inconsistencies in LLM outcomes.
Figure 1: Left: level sets of the ground truth $f^{\star}$. We perturb input $x_{0}$ to $x_{1},x_{2},\cdots$ on the isocontour so that $f^{\star}(x_{0})=f^{\star}(x_{1})=\cdots$. Right: level sets of the model $\hat{f}$. The deviations from $\hat{f}(x_{i})_{i\geq 1}$ to $\hat{f}(x_{0})$ characterize the uncertainty of $\hat{f}$.
To highlight the importance of input perturbation, we start from a proof-of-concept example where the target function $f^{\star}$ is a single output function $f^{\star}:\mathbb{R}^{d}\to\mathbb{R}$. We don’t have access to $f^{\star}$, but we have access to an approximation $\hat{f}:\mathbb{R}^{d}\to\mathbb{R}$, such as a pre-trained model. At first sight, it seems hopeless to quantify the alignment between $\hat{f}$ and $f^{\star}$ without knowing $f^{\star}$. However, the semantic equivalence class relative to an input $x_{0}$ provides a set $\mathcal{I}(x_{0})\subseteq\mathbb{R}^{d}$, such that $f^{\star}(x^{\prime})=f^{\star}(x_{0}),\,\forall x^{\prime}\in\mathcal{I}(x_{0})$. The equivalence class provides valuable information for UQ. An illustration is shown in Fig. 1.
The shape of $\mathcal{I}(x_{0})$ could be complex for general $f^{\star}$. However, in a small neighborhood of $x_{0}$ where $\nabla f^{\star}(x_{0})\neq 0$, $\mathcal{I}(x_{0})$ should be locally close to the tangent space of $f^{\star}$. More specifically, we define tangent invariance set as $\mathcal{I}_{\text{tangent}}(x_{0})=\{x_{0}+\left(I-\eta_{\nabla f^{\star}}\right)z;z\in\mathbb{R}^{d}\}$, where $\eta_{\nabla f^{\star}}=\frac{\nabla f^{\star}(x_{0})\nabla f^{\star}(x_{0})^{\top}}{\lVert\nabla f^{\star}(x_{0})\rVert^{2}}$. Taylor expansion shows $\mathcal{I}_{\text{tangent}}(x_{0})\approx\mathcal{I}(x_{0})$ when restricted to the small neighborhood of $x_{0}$. To generate algorithmic insight, we further define a probability density on $\mathcal{I}_{\text{tangent}}(x_{0})$: we use $P(x^{\prime};x_{0},\sigma)$ to denote the probability density function of $x^{\prime}=x_{0}+\left(I-\eta_{\nabla f^{\star}}\right)z$ where $z$ is sampled from a $d$-dimensional isotropic Gaussian distribution $\mathcal{N}(0,\sigma^{2})$. It is easy to verify that $P(\cdot;x_{0},\sigma)$ is a Gaussian distribution supported on $\mathcal{I}_{\text{tangent}}(x_{0})$ whose mean is $x_{0}$.
The following Lemma shows that we can estimate the angle between $\nabla\hat{f}(x)$ and $\nabla f^{\star}(x)$ by examining the variance of $\hat{f}(x^{\prime})$, where $x^{\prime}$ is sampled from $P(x^{\prime};x,\sigma)$.
Lemma 2.1
Assume (1) $\hat{f}$ is twice differentiable, and both $\|\nabla\hat{f}(x)\|$ and $\|\nabla^{2}\hat{f}(x)\|_{\texttt{op}}$ are bounded for all $x\in\mathbb{R}^{d}$, and (2) $\nabla\hat{f}(x_{0})\neq 0$ and $\nabla f^{\star}(x_{0})\neq 0$. Then, for sufficiently small $\sigma$, we have
$\frac{1}{\sigma^{2}}\mathbb{V}_{x^{\prime}\sim\mathcal{P}(\cdot;x_{0},\sigma^{2})}[\hat{f}(x^{\prime})]=\|\nabla\hat{f}(x_{0})\|^{2}\sin^{2}\theta\left(\nabla\hat{f}(x_{0}),\nabla f^{\star}(x_{0})\right)+\mathcal{O}(\sigma),$ (1)
where $\theta(v_{1},v_{2})=\arccos\left(\frac{v_{1}^{\top}v_{2}}{\|v_{1}\|\|v_{2}\|}\right)$ denotes the angle between two vectors $v_{1}$ and $v_{2}$.
Equation (1) shows that the variance of function values of $\hat{f}$ on the invariance set is indicative of the alignment between $\nabla\hat{f}$ and $\nabla f^{\star}$. When the variance is larger, $\hat{f}$ respects the invariance of $f^{\star}$ less, which in turn implies larger misalignment between $\nabla\hat{f}$ and $\nabla f^{\star}$ and larger uncertainty. The full proof of Lemma 2.1 is relegated to the Appendix B.
Despite its simplicity, Lemma 2.1 demonstrates two important elements of perturbation-based UQ: input perturbation and output variance evaluation. In the following subsections, we introduce designs that implement these elements in the LLM context to characterize the uncertainty.
2.2 A probabilistic framework via dual random walks
The variance estimate in Lemma 2.1 presents useful insights yet is oversimplified to realistically model LLMs, whose outputs are token sequences rather than a single scalar. Also, it is difficult to sample exactly from a semantic equivalence class. To tackle these challenges, we introduce a probabilistic approach inspired by Markov chains.
We use $S_{x_{0}}$ to denote a semantic input to an LLM $f$ and $S_{y_{0}}=f(S_{x_{0}})$ one of its possible corresponding outputs. Since $f$ is usually stochastic, the output $S_{y_{0}}=f(S_{x_{0}})$ is also random. Now consider a semantic embedding function $\psi$, which maps input and output samples into a continuous space $\mathcal{X}$: $(x_{0},y_{0})=(\psi(S_{x_{0}}),\psi(S_{y_{0}}))$. Also, assume we have a perturbation algorithm $\mathcal{P}er(S_{x_{0}})$ that generates a finite set of perturbed inputs:
$\mathcal{P}er(S_{x_{0}})=\{S_{x_{0}},S_{x_{1}},\dots,S_{x_{n}}\},$
where the detailed implementation of $\mathcal{P}er(S_{x_{0}})$ will be discussed in Sec. 2.5.
Applying the embedding function $\psi(S_{x_{i}})$ , we obtain the embeddings:
$X_{n}=\{x_{0},x_{1},\dots,x_{n}\},\,\text{where }x_{i}=\psi(S_{x_{i}}).$
One corresponding output embedding set is hence given as
$Y_{n}=\{y_{0},y_{1},\dots,y_{n}\},\,\text{where }y_{i}=\psi(S_{y_{i}})=\psi(f(S_{x_{i}})).$
Samples in $X_{n}$ and $Y_{n}$ are one-to-one correspondent. We can thus estimate the structural similarity of the samples in the two sets. In the following, we propose a probabilistic approach to characterize the structural similarity through the lens of a random walk.
We first introduce two Markov chains $\mathcal{M}_{x}$ and $\mathcal{M}_{y}$. They both contain the same states $\mathcal{S}=\{0,1,2,\cdots,n\}$. Assume a non-negative similarity function $a_{\text{Similarity}}(x,x^{\prime}):\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{0}^{+}$ that quantifies how close are two embeddings in $\mathcal{X}$, then we can define the transition probability matrices of $\text{P}_{x},\text{P}_{y}\in\mathbb{R}^{(n+1)\times(n+1)}$ as:
$\text{P}_{x}[i,j]$ $\triangleq\frac{a_{\text{Similarity}}(x_{i},x_{j})}{\sum_{k=0}^{n}a_{\text{Similarity}}(x_{i},x_{k})},$ (2)
$\text{P}_{y}[i,j]$ $\triangleq\frac{a_{\text{Similarity}}(y_{i},y_{j})}{\sum_{k=0}^{n}a_{\text{Similarity}}(y_{i},y_{k})}.$
The two transition probability matrices characterize two random walks in the space of $n+1$ pairs $\{x_{i},y_{i}\}_{i=0}^{n}$, where the transition probability from $i$ to $j$ is higher if their input or output semantic features are closer. Notice that $a_{\text{Similarity}}$ can be defined in different ways, some of which are explored in our numerical studies.
At its core,  (2) defines stochastic dynamics on the set $\mathcal{S}$, capturing the similarity structures in both the input and output spaces. These dual dynamics uncover meaningful semantic patterns that can be leveraged for UQ. There are various ways to characterize uncertainty by examining the alignment between the two induced graphs (Vishwanathan et al., 2010). In what follows, we construct a framework tailored for discrete input and output spaces to rigorously define uncertainty.
2.3 Constructing the distributions
Figure 2: $P(X|Y)=\text{P}_{y}\text{P}_{x}$ dynamics.
We use $X$ and $Y$ to denote discrete random variables whose supports are $X_{n}$ and $Y_{n}$, respectively. There are many possible ways to define possible distributions of $X$ and $Y$. In the following, we will introduce one design of $P(Y)$ and $P(X|Y)$ based on (2). For notational simplicity, we denote the uniform distribution over all states $\mathcal{S}$ by $\pi^{\text{Uniform}}$, which is given by $\pi^{\text{Uniform}}=\begin{bmatrix}\frac{1}{n+1}&\frac{1}{n+1}&\dots&\frac{1}{n+1}\end{bmatrix}$.
The marginal distribution-$P(Y)$
The random variable $Y$ corresponds to the LLM’s response to a question that is perturbed from the original question $x_{0}$. We define the marginal distribution of $Y$ as:
$\small\begin{split}P(Y=y_{j})\triangleq(\pi^{\text{Uniform}}\,\text{P}_{y})[j]&=\frac{1}{n+1}\sum_{i=0}^{n}\frac{a_{\text{Similarity}}(y_{i},y_{j})}{\sum_{k}a_{\text{Similarity}}(y_{i},y_{k})},\end{split}$ (3)
where notation $[j]$ denotes the $j$-th element of a vector.
The distribution (3) has an intuitive interpretation: we randomly sample a point uniformly from the state space $\mathcal{S}$, then randomly transit the sample with $\text{P}_{y}$ for one step. After the transit step, nodes whose corresponding output samples are surrounded by many similar outputs are assigned a higher probability, while isolated nodes with fewer similar neighbors are assigned a lower probability. Therefore, the mass is concentrated on regions of high semantic density in the output space.
The conditional distribution-$P(X|Y)$
Next, we introduce the conditional probability $P(X=x_{i}\mid Y=y_{j})$ as
$P(X=x_{i}|Y=y_{j})=(\text{P}_{y}\text{P}_{x})[j,i]=\sum_{k}\frac{a_{\text{Similarity}}(x_{i},x_{k})}{\sum_{m}a_{\text{Similarity}}(x_{m},x_{k})}\frac{a_{\text{Similarity}}(y_{j},y_{k})}{\sum_{l}a_{\text{Similarity}}(y_{j},y_{l})},$ (4)
where notation $[j,i]$ denotes the $(j,i)$-th entry of the matrix $\text{P}_{y}\text{P}_{x}\in\mathbb{R}^{n+1\times n+1}$.
Essentially, (4) constructs the conditional probability using the transition probability from state $i$ to $j$ under the composite dynamics. We design a two-stage process on $\mathcal{S}$. Initially, states are transitioned by $\text{P}_{y}$ that characterizes the output similarity. Then, the states are transitioned by $\text{P}_{x}$ that characterizes the input similarity. Since $\text{P}_{x}$ and $\text{P}_{y}$ are defined on the same state space $\mathcal{S}$ linked by an LLM, (4) establishes a probabilistic bridge linking similarity in the output space to the similarity in the input space. This is shown in Fig. 2.
We explicitly model the conditional distribution of input $X$ rather than output $Y$ for two reasons. First, samples generated by the perturbation algorithm $\mathcal{P}er(S_{x_{0}})$ may deviate from $x_{0}$ in meanings. We thus reweight these samples based on their pairwise similarity in the input space and output space. Second, we observe there exists an inherent asymmetry in semantic variability between inputs and outputs in LLMs: multiple semantically distinct prompts can often elicit very similar outputs, whereas small changes to an input typically result in only modest variation in output. As a result, modeling the conditional distribution of $X$ given $Y$ provides a more stable and reliable basis for inferring the diversity of inputs that are semantically related to an output.
Our goal is to leverage the information from $Y$ to guide the conditional distribution of $X$. This interplay between input and output similarities allows the probability distribution to adapt to the structure of the questions and responses, ensuring that inputs are assigned higher probability when they are supported by multiple semantically similar input-output pairs.
The distributions derived from $P(Y)$ and $P(X|Y)$
Based on defined $P(Y)$ and $P(X|Y)$, we can calculate the joint probability $P(X=x_{i},Y=y_{j})\triangleq P(x_{i}|y_{j})P(y_{j})$, and the marginal probability $P(X=x_{i})\triangleq\sum_{j=0}^{n}P(x_{i}|y_{j})P(y_{j})=\pi^{\text{Uniform}}\text{P}_{y}\text{P}_{y}\text{P}_{x}[i]$. We are also ready to apply Bayes’ theorem to compute the conditional probability $P(Y=y_{j}|X=x_{i})\triangleq P(y_{j}|x_{i})=\frac{P(x_{i}|y_{j})P(y_{j})}{\sum_{k=0}^{n}P(x_{i}|y_{k})P(y_{k})}$.
Building on this framework, researchers can flexibly define uncertainty measures tailored to their needs. One of our main contributions is to offer this general probabilistic foundation based on $P(Y)$ and $P(Y|X)$, which supports a wide range of definitions such as divergence-based metrics (for example, KL divergence and Wasserstein distance), entropy-based measures derived from various distributions, and other probabilistic constructs. As an example, we introduce an entropy-based metric in the following section.
2.4 Inv-Entropy via bootstrapping and Monte Carlo
We next introduce how to leverage the probabilistic framework described above to define our UQ measure, denoted as Inverse-Entropy (Inv-Entropy).
Inv-Entropy
A natural measure of uncertainty is the conditional sample entropy $H(X_{n}\mid Y_{n})$, which quantifies the ambiguity of the input set $X_{n}$ given the corresponding output set $Y_{n}$:
$H(X_{n}\mid Y_{n})\triangleq-\operatorname{trace}(\text{P}_{y}\text{P}_{x}\odot\log(\text{P}_{y}\text{P}_{x}))=-\sum_{i=0}^{n}P(x_{i}\mid y_{i})\log P(x_{i}\mid y_{i}),$ (5)
where $\odot$ denotes a Hadamard product. This quantity reflects the alignment between $\text{P}_{y}$ and $\text{P}_{x}$. High entropy indicates that semantically similar inputs lead to different output or semantically different outputs lead to similar input, revealing uncertainty in the model’s behavior.
That said, it is important to realize that $X_{n}$ does not map to a unique $Y_{n}$ due to the inherent stochasticity of $f$, as LLMs can produce different outputs for the same input. Hence, beyond perturbations that capture epistemic uncertainty, replications are needed to capture aleatoric uncertainty from sampling randomness. To this end, we replicate multiple responses per question: each perturbation $S_{x_{i}}\in\mathcal{P}er(S_{x_{0}})$ is passed to the LLM $r$ times, generating a replicated set $\mathcal{R}_{i}=\{y_{i,1},\dots,y_{i,r}\}$, where $y_{i,\cdot}=\psi(f(S_{x_{i}}))$.
Bootstrapping & Monte Carlo
With this, we are now able to generate $S$ bootstrapped samples $Y_{n}^{(s)}=\{y_{0}^{(s)},\dots,y_{n}^{(s)}\}$ for $s\in\{1,\dots,S\}$ corresponding to $X_{n}$, where $y_{i}^{(s)}\sim\mathcal{R}_{i}$ for $i\in\{0,\dots,n\}$ is sampled with replacement. Each bootstrapped output embedding set $Y_{n}^{(s)}$, together with the input embedding set $X_{n}$, induces a probabilistic framework that yields an instance of our UQ measure, Inv-Entropy. Accordingly, we apply Monte Carlo estimation to compute the average Inv-Entropy across these instances, resulting in the final UQ measure shown below:
$\widehat{H}(X\mid Y)=\frac{1}{S}\sum_{s=1}^{S}H(X_{n}\mid Y_{n}^{(s)}).$ (6)
Our overall framework in given in Algorithm 1. It is important to emphasize that a key strength of our framework lies in its flexibility: it accommodates arbitrary choices of embedding function $\psi$, perturbation strategy $\mathcal{P}er$, and similarity metric $a_{\text{Similarity}}$, allowing it to adapt to different tasks and model architectures. Moreover, $\widehat{H}(X\mid Y)$ represents just one possible uncertainty measure; many others can be defined within our probabilistic framework. For instance, our numerical studies evaluate alternatives such as the Wasserstein distance between the marginal distributions $P(X)$ and $P(Y)$.
Algorithm 1 Inv-Entropy overall framework
1:Input: $(S_{x_{0}},S_{y_{0}})$, $f$, $\psi$, $\mathcal{P}er$, and $a_{\text{Similarity}}$
2:Perturb: Use $\mathcal{P}er(S_{x_{0}})$ and $\psi$ to obtain $X_{n}$, Compute $\text{P}_{x}$ using (2)
3:LLM Generation: Input each element in $\mathcal{P}er(S_{x_{0}})$, $r$ times into $f$ and obtain $\{\mathcal{R}_{i}\}_{i=0}^{n}$.
4:for $s=1$ to $S$ do
5:     Sample $Y_{n}^{(s)}=\{y_{0}^{(s)},\dots,y_{n}^{(s)}\}$
6:     Compute $\text{P}_{y}^{(s)}$ using (2)
7:     Compute $P^{(s)}(x_{i}|y_{i})=\text{P}^{(s)}_{y}\cdot\text{P}_{x}[i,i]$ (from (4)) for $\forall$ $i$
8:     Compute $H(X_{n}\mid Y_{n}^{(s)})$ using (5)
9:end for
10:Compute $\widehat{H}(X|Y)$ using (6)
2.5 GAAP
In this section, we present a genetic algorithm-based adversarial perturbation (GAAP) that progressively modifies the semantic input $S_{x_{0}}$ to generate controlled perturbations, $\mathcal{P}er(S_{x_{0}})$. Below we highlight our overarching framework, while algorithmic details are relegated to Appendix C. As shown in Fig. 3, the process consists of an initialization step and multiple iterative procedures. In the initialization step, we construct a population $\text{Pop}_{0}(S_{x_{0}})$ consisting of perturbed versions of $S_{x_{0}}$. More specifically, each text in $\text{Pop}_{0}(S_{x_{0}})$ is derived from $S_{x_{0}}$ by replacing one keyword with a synonym, hypernym, hyponym from WordNet (Miller, 1995), or a deletion.
Figure 3: Illustration of GAAP on a TriviaQA (Joshi et al., 2017) question.
With an initial $\text{Pop}_{0}(S_{x_{0}})$, GAAP updates the $\text{Pop}_{t}(S_{x_{0}})$ through subsequent steps of crossovers and mutations. In the crossover step, we first select a random subset of $\text{Pop}_{t}(S_{x_{0}})$ based on $a_{\text{Similarity}(x_{0},\cdot)}$, such that samples are chosen with higher probability if they are closer to $x_{0}$. Next, we randomly segment each selected sentence. These sentence segments are then randomly concatenated to generate new sentences, as illustrated in Crossover of Fig. 3.
In the mutation step, we perturb the recombined sentences with further key word substitutions and deletions, which introduce additional variations to the cross-pollinated texts. We construct the next generation population $\text{Pop}_{t+1}(S_{x_{0}})$ as the union of the selected, crossovered, and mutated texts.
GAAP proceeds by iteratively updating $\text{Pop}_{t}(S_{x_{0}})$ with crossovers and mutations for $T$ iterations or until all texts in $\text{Pop}_{t}(S_{x_{0}})$ have similarity to $S_{x_{0}}$ smaller than a predefined constant $\delta$. Finally, we construct $\mathcal{P}er(S_{x_{0}})$ by sampling from populations at different generations $\{\text{Pop}_{t}(S_{x_{0}})\}_{t=0,\tau,2\tau,\cdots}$, where $\tau\in\mathbb{Z}$ is a fixed gap. Since texts in $\text{Pop}_{t}(S_{x_{0}})$ tend to deviate from $S_{x_{0}}$ further with larger $t$, such construction of $\mathcal{P}er(S_{x_{0}})$ ensures diverse representation of perturbed texts with different levels of similarity to $S_{x_{0}}$.
3 Experiments
Table 1: Comparison of AUROC, PRR, and Brier scores across all the 5 datasets. We use GPT-3.5-Turbo with ChatGPT-based paraphrasing, and DeBERTa-v2-xlarge-MNLI embedding function. Bold and underline denote the best and second-best performers, respectively.
Metric Method Datasets
TriviaQA SciQ NQ MMLU GSM8K
AUROC ($\uparrow$) Semantic Entropy 0.579 $\pm$ 0.044 0.679 $\pm$ 0.045 0.661 $\pm$ 0.048 0.518 $\pm$ 0.048 0.589 $\pm$ 0.052
VU 0.695 $\pm$ 0.060 0.480 $\pm$ 0.060 0.549 $\pm$ 0.057 0.523 $\pm$ 0.054 0.557 $\pm$ 0.057
P(True) 0.604 $\pm$ 0.050 0.522 $\pm$ 0.026 0.559 $\pm$ 0.033 0.474 $\pm$ 0.027 0.571 $\pm$ 0.056
LexSim 0.649 $\pm$ 0.055 0.681 $\pm$ 0.046 0.653 $\pm$ 0.053 0.643 $\pm$ 0.054 0.598 $\pm$ 0.060
DegMat 0.734 $\pm$ 0.056 0.672 $\pm$ 0.059 0.616 $\pm$ 0.057 0.608 $\pm$ 0.058 0.678 $\pm$ 0.059
LUQ 0.637 $\pm$ 0.067 0.726 $\pm$ 0.048 0.675 $\pm$ 0.058 0.648 $\pm$ 0.057 0.662 $\pm$ 0.064
KLE 0.333 $\pm$ 0.054 0.341 $\pm$ 0.056 0.357 $\pm$ 0.053 0.360 $\pm$ 0.064 0.338 $\pm$ 0.061
Inv-Entropy 0.788 $\pm$ 0.054 0.740 $\pm$ 0.050 0.703 $\pm$ 0.060 0.780 $\pm$ 0.041 0.695 $\pm$ 0.051
NI-Entropy 0.786 $\pm$ 0.057 0.681 $\pm$ 0.056 0.669 $\pm$ 0.057 0.710 $\pm$ 0.052 0.650 $\pm$ 0.069
NR-Inv-Entropy 0.743 $\pm$ 0.061 0.720 $\pm$ 0.049 0.659 $\pm$ 0.063 0.604 $\pm$ 0.059 0.677 $\pm$ 0.064
WD-px-py 0.518 $\pm$ 0.060 0.303 $\pm$ 0.060 0.443 $\pm$ 0.056 0.573 $\pm$ 0.061 0.605 $\pm$ 0.069
MAX-py-x 0.723 $\pm$ 0.054 0.674 $\pm$ 0.054 0.608 $\pm$ 0.053 0.585 $\pm$ 0.059 0.618 $\pm$ 0.059
PRR ($\uparrow$) Semantic Entropy 0.517 $\pm$ 0.060 0.763 $\pm$ 0.044 0.673 $\pm$ 0.054 0.690 $\pm$ 0.058 0.335 $\pm$ 0.056
VU 0.723 $\pm$ 0.053 0.677 $\pm$ 0.053 0.584 $\pm$ 0.060 0.654 $\pm$ 0.055 0.328 $\pm$ 0.057
P(True) 0.797 $\pm$ 0.042 0.679 $\pm$ 0.050 0.599 $\pm$ 0.049 0.671 $\pm$ 0.041 0.303 $\pm$ 0.056
LexSim 0.810 $\pm$ 0.045 0.770 $\pm$ 0.051 0.654 $\pm$ 0.062 0.767 $\pm$ 0.053 0.356 $\pm$ 0.076
DegMat 0.882 $\pm$ 0.041 0.802 $\pm$ 0.046 0.644 $\pm$ 0.066 0.771 $\pm$ 0.058 0.462 $\pm$ 0.091
LUQ 0.854 $\pm$ 0.043 0.840 $\pm$ 0.045 0.678 $\pm$ 0.066 0.787 $\pm$ 0.052 0.504 $\pm$ 0.094
KLE 0.704 $\pm$ 0.048 0.592 $\pm$ 0.059 0.495 $\pm$ 0.056 0.612 $\pm$ 0.061 0.224 $\pm$ 0.043
Inv-Entropy 0.885 $\pm$ 0.044 0.853 $\pm$ 0.042 0.764 $\pm$ 0.064 0.898 $\pm$ 0.030 0.521 $\pm$ 0.094
NI-Entropy 0.883 $\pm$ 0.043 0.781 $\pm$ 0.053 0.689 $\pm$ 0.071 0.823 $\pm$ 0.055 0.501 $\pm$ 0.098
NR-Inv-Entropy 0.840 $\pm$ 0.054 0.844 $\pm$ 0.045 0.709 $\pm$ 0.068 0.743 $\pm$ 0.064 0.518 $\pm$ 0.087
WD-px-py 0.763 $\pm$ 0.051 0.587 $\pm$ 0.056 0.575 $\pm$ 0.060 0.777 $\pm$ 0.054 0.420 $\pm$ 0.085
MAX-py-x 0.875 $\pm$ 0.038 0.821 $\pm$ 0.048 0.712 $\pm$ 0.055 0.749 $\pm$ 0.062 0.413 $\pm$ 0.081
Brier ($\downarrow$) Semantic Entropy 0.166 $\pm$ 0.023 0.173 $\pm$ 0.020 0.208 $\pm$ 0.017 0.208 $\pm$ 0.020 0.188 $\pm$ 0.018
VU 0.160 $\pm$ 0.022 0.196 $\pm$ 0.017 0.234 $\pm$ 0.010 0.219 $\pm$ 0.017 0.188 $\pm$ 0.020
P(True) 0.172 $\pm$ 0.022 0.215 $\pm$ 0.017 0.233 $\pm$ 0.012 0.215 $\pm$ 0.015 0.189 $\pm$ 0.021
LexSim 0.151 $\pm$ 0.024 0.179 $\pm$ 0.020 0.202 $\pm$ 0.016 0.187 $\pm$ 0.020 0.174 $\pm$ 0.019
DegMat 0.140 $\pm$ 0.021 0.164 $\pm$ 0.018 0.213 $\pm$ 0.014 0.191 $\pm$ 0.018 0.156 $\pm$ 0.019
LUQ 0.148 $\pm$ 0.020 0.159 $\pm$ 0.016 0.191 $\pm$ 0.019 0.180 $\pm$ 0.019 0.151 $\pm$ 0.019
KLE 0.188 $\pm$ 0.021 0.218 $\pm$ 0.016 0.241 $\pm$ 0.006 0.213 $\pm$ 0.018 0.193 $\pm$ 0.021
Inv-Entropy 0.128 $\pm$ 0.020 0.157 $\pm$ 0.018 0.182 $\pm$ 0.020 0.147 $\pm$ 0.017 0.152 $\pm$ 0.020
NI-Entropy 0.124 $\pm$ 0.020 0.164 $\pm$ 0.017 0.196 $\pm$ 0.018 0.168 $\pm$ 0.021 0.156 $\pm$ 0.022
NR-Inv-Entropy 0.138 $\pm$ 0.021 0.159 $\pm$ 0.015 0.199 $\pm$ 0.019 0.188 $\pm$ 0.021 0.165 $\pm$ 0.021
WD-px-py 0.184 $\pm$ 0.019 0.212 $\pm$ 0.016 0.232 $\pm$ 0.009 0.188 $\pm$ 0.018 0.169 $\pm$ 0.021
MAX-py-x 0.148 $\pm$ 0.019 0.177 $\pm$ 0.017 0.207 $\pm$ 0.012 0.189 $\pm$ 0.020 0.169 $\pm$ 0.018
Models and tasks
We conducted experiments using two language models: GPT-3.5-Turbo, a black-box model accessed via API, and LLaMA-3.1-8B-Instruct, a grey-box model. We evaluated our framework on datasets spanning three categories: question answering (TriviaQA (Joshi et al., 2017), SciQ (Welbl et al., 2017), Natural Questions (Kwiatkowski et al., 2019)), multiple choice (MMLU (Hendrycks et al., 2020)), and mathematical reasoning (GSM8K (Cobbe et al., 2021)).
Baselines
We compared our method with various state-of-the-art benchmarks highlighted in Sec. 1, Appendix A and a recent paper (Vashurin et al., 2024) identifying them as top-performers. These include: Semantic Entropy (Farquhar et al., 2024), Verbalized Uncertainty (VU)(Tian et al., 2023), P(True)(Kadavath et al., 2022), Lexical Similarity (LexSim)(Fomicheva et al., 2020), Degree Matrix (DegMat)(Lin et al., 2024), Long-text Uncertainty Quantification (LUQ)(Zhang et al., 2024a), Kernel Language Entropy (KLE)(Nikitin et al., 2024). We also include additional UQ measures based on our framework: (i) NI-Entropy: Non-inverse entropy which uses $P(Y\mid X)$ derived in Sec. 2.3 instead of $P(X\mid Y)$; the rest remains the same. (ii) NR-Inv-Entropy: entropy in (5) without replications. (iii) WD-px-py: Wasserstein distance $\mathsf{WD}(P(X),P(Y))$ ; (iv) MAX-py-x: $\max_{i}P(y_{i}|x_{i})$ from (3).
Evaluation metrics
We evaluate benchmarks using four metrics that are grouped into calibration and uncertainty-based categories. Calibration-based metrics, AUROC, PRR (Malinin and Gales, 2021), and Brier Score (Brier, 1950), assess how well confidence aligns with correctness. The correctness is calculated in the following way. On MMLU, we use exact match of results. On other datasets, we use GPT-3.5-Turbo to judge semantic equivalence. Confidence is defined as the negative of the uncertainty measure for AUROC and PRR, while for Brier Score, we apply isotonic normalization (Zadrozny and Elkan, 2002) to map uncertainty measures to the [0,1] range for all methods. Using correctness as a proxy for UQ has its limitations in open-ended applications or weakly supervised settings where correctness information is not available or not well defined. To circumvent the dependence on correctness, we introduce a new evaluation metric called Temperature Sensitivity of Uncertainty (TSU), a metric that quantifies the proportion of instances where uncertainty increases with temperature. TSU is inspired by the property that LLMs generate responses using a temperature-scaled softmax function, which controls the randomness of token sampling. Higher temperatures flatten token distribution, increasing randomness and uncertainty, while lower temperatures make the sampling process more deterministic (Hinton et al., 2015). Formally, given a sequence of temperature values $\mathbb{T}_{1}<\mathbb{T}_{2}<\dots<\mathbb{T}_{n}$, TSU is defined as:
$\text{TSU}(\mathbb{T}_{1},\mathbb{T}_{2},\dots,\mathbb{T}_{n})=\frac{1}{|\mathcal{D}|}\sum_{S_{x}\in\mathcal{D}}\mathbb{I}\Big{(}\texttt{UQ}(S_{x},\mathbb{T}_{1})<\texttt{UQ}(S_{x},\mathbb{T}_{2})<\dots<\texttt{UQ}(S_{x},\mathbb{T}_{n})\Big{)},$ (7)
where $\mathcal{D}$ is the dataset, $S_{x}$ is a question in this dataset, $\texttt{UQ}(S_{x},\mathbb{T})$ represents a UQ subroutine (such as Inv-Entropy) for input $S_{x}$ at temperature $\mathbb{T}$, and $\mathbb{I}(\cdot)$ is the indicator function, which equals 1 if the condition holds and 0 otherwise. A salient feature in the definition of TSU in (7) is that it only depends on $S_{x}$, thus is agnostic to the “ground truth” output $y$. In addition, TSU extends beyond conventional calibration metrics by evaluating the granularity of uncertainty estimation. By leveraging temperature scaling as a probing mechanism, TSU assesses how effectively a method distinguishes between gradations of uncertainty.
Implementation details
Our framework requires three inputs: $\psi$, $\mathcal{P}er$, and $a_{\text{Similarity}}$. For $\mathcal{P}er$, we apply two strategies: (1) ChatGPT-based paraphrasing, generating nine perturbed versions per question, (2) GAAP introduced in Sec. 2.5 with a similarity threshold of $\delta=0.7$. For $\psi$, we employ three state-of-the-art approaches: (i) SBERT-small (paraphrase-MiniLM-L6-v2), (ii) SBERT-large (all-mpnet-base-v2) (Reimers and Gurevych, 2019), and (iii) DeBERTa-v2-xlarge-MNLI (He et al., 2021). For (i) and (ii), we use cosine similarity $a_{\text{Similarity}}(x,x^{\prime})=(1+\cos(x,x^{\prime}))/2$. While, (iii) generates an entailment score; however, this score is not symmetric. To address this, we take the average $(a_{\text{Similarity}}(x,x^{\prime})+a_{\text{Similarity}}(x^{\prime},x))/2$. All experiments were conducted with an NVIDIA A100 GPU. Detailed experimental set up including prompts and parameters used are detailed in Appendix D. Appendix D also includes additional simulation results; we present only the core findings in the main paper for clarity and focus.
3.1 Results
Calibration-based
As shown in Table 1, our proposed Inv-Entropy method demonstrates consistently strong and stable performance across all five datasets and metrics. While it does not always rank first, it frequently appears among the top performers and achieves notable results, such as 0.788 AUROC on TriviaQA and 0.898 PRR on MMLU. It also ranks among the top two in Brier score, indicating well-calibrated confidence estimates. Its variant, NR-Inv-Entropy, also performs competitively, underscoring the robustness of our approach and the added benefit of incorporating replications. We intentionally use ChatGPT-based paraphrasing to highlight the advantages of our method independent of GAAP. Results for LLaMA-3.1-8B-Instruct are deferred to Appendix D.
TSU
Table 2 reports TSU results across all five datasets. These results align with the calibration-based UQ metrics in Table 1: methods defined by our probabilistic framework consistently achieve top performance, while LexSim, DegMat and LUQ exhibit highly inconsistent results, occasionally leading in specific cases but underperforming in others. Methods like P(True) perform poorly in TSU due to their binary nature, which fundamentally lacks granularity. Similarly, Semantic Entropy shows limited discriminatory power as its values often cluster around few discrete points. In contrast to these coarse-grained approaches, INV-Entropy is fine-grained with superior ability to reflect uncertainty variations, demonstrated by consistent top TSU performance across all datasets and temperature settings. This result highlights the probabilistic framework’s ability to capture uncertainty trends beyond labeled datasets.
Performance breakdown
(i) Impact of replications: The comparison between NR-Inv-Entropy and Inv-Entropy in Tables 1 and 2 highlights the impact of replication and bootstrapping. Notably, NR-Inv-Entropy often ranks second, showcasing the strength of our framework even without replication. This suggests that one can trade off a small loss in accuracy for fewer queries. Appendix D includes an ablation over $S$ and $r$, further confirming this finding. (ii) Impact of inversion: The comparison between NI-Entropy and Inv-Entropy further confirms the advantages of our inverse approach, which examines the diversity of inputs that could have led to a specific output. (iii) Framework generality: The often competitive performance of WD-px-py and MAX-py-x highlights the robustness of our framework and its flexibility in defining a wide range of UQ measures. (iv) Impact of perturbations: Fig. 4 shows that GAAP consistently improves AUROC across all three datasets and both LLMs compared to ChatGPT-based paraphrasing. By principledly enhancing input diversity, GAAP better tests model robustness and improves uncertainty quantification. These results underscore the importance of meaningful perturbations. (v) Impact of embedding function: Our framework delivers consistently strong uncertainty estimates with every encoder we tested, including SBERT (paraphrase-MiniLM-L6-v2), SBERT (all-mpnet-base-v2), and DeBERTa. Although the larger encoders provide slightly higher scores on several tasks, the overall gap is modest, showing that even lightweight models can support reliable UQ when paired with our method. The close alignment between results from SBERT (all-mpnet-base-v2) and DeBERTa further suggests that entailment and similarity signals extracted by the two architectures contain overlapping information.
Figure 4: AUROC of Inv-Entropy under different perturbation methods (GAAP or ChatGPT-based paraphrasing) and embedding functions, on both ChatGPT and LLaMA models.
[…]
Method
TSU (0.3, 0.7)
TSU (0.7, 1.0)
TSU (1.0, 1.4)
TSU (0.3, 1.0)
TSU (0.3, 1.4)
TSU (0.7, 1.4)
TSU (0.3-1.0)
TSU (0.7-1.4)
TSU (0.3-1.4)
TriviaQA
Semantic Entropy 11.57 14.26 17.35 21.56 25.51 17.35 5.18 5.35 3.94
VU 22.45 27.55 38.78 25.51 41.84 42.86 0.00 4.92 0.00
P(True) 1.02 2.13 3.85 1.38 0.98 3.49 0.00 0.00 0.00
LexSim 22.39 22.95 46.94 30.61 53.36 53.54 9.18 12.38 8.16
DegMat 24.58 33.21 45.37 31.77 48.98 47.96 18.37 20.02 13.27
LUQ 20.41 31.08 48.06 33.67 52.34 50.00 14.78 27.63 10.20
KLE 7.14 17.35 13.45 4.57 1.28 6.42 2.79 1.31 0.00
Inv-Entropy 52.42 66.33 77.55 76.53 92.86 88.78 30.49 47.21 19.05
NI-Entropy 55.14 47.96 61.22 59.35 67.35 60.20 21.43 19.32 11.73
WD-px-py 44.81 67.35 57.62 60.20 64.49 69.39 23.11 36.73 12.06
MAX-py-x 65.31 50.83 74.49 67.35 85.71 81.63 27.42 32.65 16.03
SciQ
Semantic Entropy 11.09 15.48 21.55 17.35 27.62 25.51 6.73 4.31 2.65
VU 21.43 33.67 28.57 35.71 31.42 32.65 1.55 6.12 0.00
P(True) 0.00 3.42 3.14 0.00 1.33 0.79 0.00 0.00 0.00
LexSim 31.73 29.59 47.96 40.82 61.22 59.18 19.39 13.27 11.22
DegMat 39.80 47.96 57.14 52.04 64.28 64.46 23.47 23.99 13.04
LUQ 38.45 43.88 41.84 44.90 57.14 44.95 18.37 17.35 6.12
KLE 6.12 39.54 19.32 4.98 5.67 18.37 2.30 5.10 0.00
Inv-Entropy 46.70 62.24 72.38 66.72 75.84 80.61 22.45 35.71 14.31
NI-Entropy 42.30 59.18 55.19 48.98 54.77 63.27 21.43 22.65 9.18
WD-px-py 50.00 59.16 63.25 59.37 71.40 69.39 22.13 29.59 11.22
MAX-py-x 51.02 61.22 67.35 61.22 69.86 73.25 29.59 34.69 15.29
NQ
Semantic Entropy 22.53 20.41 35.71 27.55 39.80 30.78 4.08 10.20 2.04
VU 23.71 22.68 40.21 21.65 36.08 36.08 4.12 3.09 0.00
P(True) 4.08 3.06 2.04 1.02 2.04 1.02 0.00 0.00 0.00
LexSim 51.55 54.64 59.79 63.73 79.38 74.23 36.08 32.99 17.44
DegMat 51.04 51.04 59.38 65.62 76.04 71.88 32.29 33.33 21.42
LUQ 43.16 49.47 67.37 57.89 76.84 74.74 24.21 30.41 13.68
KLE 6.38 28.72 21.28 6.38 4.26 12.77 2.13 4.26 0.00
Inv-Entropy 61.86 59.79 78.35 74.23 91.75 85.57 30.93 41.24 22.43
NI-Entropy 47.42 60.82 59.79 58.76 70.10 70.10 23.71 29.81 10.31
WD-px-py 63.55 50.52 70.10 67.01 76.29 72.16 27.84 25.77 16.92
MAX-py-x 55.67 62.89 73.20 66.30 84.54 82.47 28.87 39.18 18.56
MMLU
Semantic Entropy 21.43 17.35 33.20 21.83 42.08 39.80 7.14 4.93 2.09
VU 23.56 19.01 37.62 30.18 32.50 38.73 1.37 2.59 0.00
P(True) 1.92 4.56 5.87 4.92 5.02 5.79 0.00 0.00 0.00
LexSim 33.94 41.17 55.06 50.00 68.37 61.22 24.78 30.61 15.28
DegMat 41.84 54.08 69.39 53.76 78.57 77.55 21.46 32.58 14.34
LUQ 53.46 47.96 61.22 58.92 68.37 62.24 27.55 27.55 10.80
KLE 10.20 25.51 26.53 7.14 4.76 12.23 2.93 2.67 0.00
Inv-Entropy 60.08 67.35 73.47 79.59 90.82 86.73 34.31 43.88 18.37
NI-Entropy 56.12 54.63 50.00 67.35 52.45 59.38 19.39 18.37 7.14
WD-px-py 50.00 59.16 66.38 59.37 70.41 69.39 25.60 29.59 11.22
MAX-py-x 62.24 62.35 65.41 73.47 81.73 80.61 31.62 25.51 14.42
GSM8K
Semantic Entropy 44.90 56.12 35.71 71.43 77.55 62.24 20.41 13.27 4.08
VU 11.34 39.18 29.90 35.05 35.05 38.14 2.06 6.19 1.03
P(True) 5.10 17.35 6.12 3.32 3.98 11.22 0.00 0.00 0.00
LexSim 54.17 63.54 54.17 65.62 64.58 63.54 30.31 23.96 10.42
DegMat 55.79 64.21 55.79 72.63 75.79 70.53 29.47 29.47 10.53
LUQ 64.95 74.23 72.16 83.51 89.69 81.44 44.33 54.64 37.11
KLE 17.02 35.11 27.66 4.23 3.72 23.40 0.00 3.19 0.00
Inv-Entropy 73.68 67.33 68.42 86.32 93.68 72.63 45.81 44.21 30.53
NI-Entropy 53.61 58.76 51.55 61.16 57.73 59.79 25.77 20.62 6.19
WD-px-py 54.08 53.06 68.37 55.10 64.29 60.20 20.41 27.43 7.14
MAX-py-x 71.88 65.62 62.50 89.58 91.67 77.08 42.71 37.50 28.12
4 Conclusion
We present a fully probabilistic framework for uncertainty quantification that models the conditional distribution of inputs given outputs through a dual random walk formulation. This inverse modeling perspective enables a principled characterization of uncertainty by capturing the semantic diversity of inputs associated with a given output. A key strength of our framework is its flexibility, allowing researchers to freely combine embedding functions, perturbation strategies, and similarity metrics to define customized uncertainty measures. As an instantiation of this idea, we introduce Inv-Entropy, a novel uncertainty metric derived from the framework. Together with the proposed perturbation algorithm GAAP and evaluation metric TSU, our method achieves state-of-the-art performance across multiple datasets. We believe this framework opens up broad opportunities for future research, providing a general foundation upon which new uncertainty measures, tailored to different purposes, can be systematically developed. We acknowledge that, like other perturbation and replication-based UQ methods, our approach may face practical limitations due to computational cost. A promising direction is to adaptively determine when further perturbation is unnecessary. We hope GAAP’s sequential design can provide a step towards this goal.
Appendix A Related Work
UQ in LLMs has been gaining increasing interest recently [Huang et al., 2024]. The existing, albeit limited, literature can generally be categorized into three perspectives.
Self-evaluation-based UQ:
Self-evaluation techniques for LLMs [Chen et al., 2024] and the verbal expression of uncertainty [Lin et al., 2022] have recently been explored to enhance interpretability and reliability. For instance, recent approaches for evaluating free-form generation tasks frequently utilize LLMs as evaluators [Zheng et al., 2023]. Additionally, for gray-box models, where the internal workings are partially known, perplexity [Chen et al., 1998] and entropy [Shannon, 1948] can be directly computed from the output logits, providing a natural UQ measure.
Replication-based UQ:
These methods generate multiple outputs for a given input and measure the deviation between them to estimate uncertainty [Grewal et al., 2024, Wagner et al., 2024, Kuhn et al., 2023]. Perhaps most prevalent is semantic entropy [Farquhar et al., 2024], which computes uncertainty by clustering semantically equivalent answers from multiple responses and calculating the entropy of the resulting clusters. While effective in capturing aleatoric uncertainty, these approaches struggle with confidently wrong predictions, as resampling often yields similar incorrect results, leading to overconfidence and poor calibration. This issue exacerbates the challenges of handling hallucinations in LLMs [Grewal et al., 2024].
Perturbation-based UQ:
This is a more recent approach that involves systematic perturbations of inputs or latent representations to evaluate output variability [Zhang et al., 2023]. While Dong et al. [2023] systematically evaluates LLM robustness for noisy slot-filling tasks under diverse input perturbations, SPUQ [Gao et al., 2024] provides a UQ metric by analyzing response variations to perturbed inputs. Notably, SPUQ achieves significant improvements in model uncertainty calibration, reducing Expected Calibration Error (ECE) by an average of 50%. Indeed, our theoretical argument in Sec. 2.1 provides a theoretical justification for the need for perturbation in effective UQ.
In light of existing literature, our method unifies replication-based and perturbation-based UQ while introducing a Bayesian perspective to LLM uncertainty estimation by modeling the posterior distribution of inputs conditioned on outputs. This Bayesian inverse design approach provides a new framework and perspective for UQ of semantic models.
Appendix B Proof of Lemma 2.1
In this section, we present the proof of Lemma 2.1. the proof is based on Taylor expansion of $\hat{f}(x^{\prime})$.
The order-2 Taylor expansion for $\hat{f}(x_{0}+(I-\eta_{\nabla f^{\star}})z)$ is,
$\hat{f}(x_{0}+(I-\eta_{\nabla f^{\star}})z)=\hat{f}(x_{0})+\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})z+R(z),$ (8)
where $R(z)$ is the remainder term defined as,
$R(z)=\frac{1}{2}z^{\top}(I-\eta_{\nabla f^{\star}})\nabla^{2}\hat{f}(x_{0}+\xi(z)(I-\eta_{\nabla f^{\star}})z)(I-\eta_{\nabla f^{\star}})z,$ (9)
where $\xi(z)\in[0,1]$ is a constant dependent on $z$.
By assumption, there exists a constant $G>0$ such that $\lVert\nabla^{2}f(x)\rVert_{op}\leq G$, $\forall x\in\mathbb{R}^{d}$. Therefore,
$R(z)\leq G\lVert(I-\eta_{\nabla f^{\star}})z\rVert^{2}/2\leq G\lVert z\rVert^{2}/2$ (10)
Since $z$ admits multivariate Gaussian distribution, we can provide an upper bound for $\mathbb{E}(R(z))$ and $\mathbb{E}(R^{2}(z))$ as
$\mathbb{E}(R(z))\leq Gd\sigma^{2}/2,$ (11)
and
$\mathbb{E}(R(z)^{2})\leq\frac{G^{2}}{4}\mathbb{E}[\lVert z\rVert^{4}]$ (12)
$\leq\frac{G^{2}}{4}d^{2}\mathbb{E}[z_{i}^{4}]=\frac{G^{2}}{4}d^{2}3\sigma^{4}.$
We can also calculate the expectation of $\hat{f}(x_{0}+(I-\eta_{\nabla f^{\star}})z)$ as
$\mathbb{E}[\hat{f}(x_{0}+(I-\eta_{\nabla f^{\star}})z)]=\hat{f}(x_{0})+\mathbb{E}(R(z)).$ (13)
Then, the variance of $\hat{f}(x_{0}+(I-\eta_{\nabla f^{\star}})z)$ is,
$\mathbb{V}[\hat{f}(x_{0}+(I-\eta_{\nabla f^{\star}})z)]=\mathbb{E}\left[\left(\hat{f}(x_{0}+(I-\eta_{\nabla f^{\star}})z)-\mathbb{E}[\hat{f}(x_{0}+(I-\eta_{\nabla f^{\star}})z)]\right)^{2}\right]$ (14)
$=\mathbb{E}\left[\left(\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})z\right)^{2}\right]+\mathbb{E}\left[\left(\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})z\right)\zeta(z)\right]+\mathbb{E}\left[\zeta(z)^{2}\right],$
where $\zeta(z)$ is defined as
$\zeta(z)=R(z)-\mathbb{E}(R(z)).$ (15)
Notice that the first term in (14) is,
$\mathbb{E}\left[\left(\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})z\right)^{2}\right]$ (16)
$=\mathbb{E}\left[\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})zz^{\top}(I-\eta_{\nabla f^{\star}})\nabla\hat{f}(x_{0})\right]$
$=\sigma^{2}\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})\nabla\hat{f}(x_{0})$
$=\sigma^{2}\left(\lVert\nabla\hat{f}(x_{0})\rVert^{2}-\frac{\left(\nabla f^{\star}(x_{0})^{\top}\nabla\hat{f}(x_{0})\right)^{2}}{\lVert\nabla f^{\star}(x_{0})\rVert^{2}}\right)$
$=\sigma^{2}\lVert\nabla\hat{f}(x_{0})\rVert^{2}\sin^{2}\theta(\nabla f^{\star}(x_{0}),\nabla\hat{f}(x_{0})).$
The third term in (14) is upper bounded by
$\mathbb{E}\left[\zeta(z)^{2}\right]$ (17)
$\leq 2\mathbb{E}[R(z)^{2}]+2\mathbb{E}[R(z)]^{2}$
$\leq 2G^{2}d^{2}\sigma^{4}=O(\sigma^{4}),$
where we used (11) and (12) in the third inequality.
The second term in (14) is upper bounded by the Holder inequality,
$\mathbb{E}\left[\left(\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})z\right)\zeta(z)\right]$ (18)
$\leq\sqrt{\mathbb{E}\left[\left(\nabla\hat{f}(x_{0})^{\top}(I-\eta_{\nabla f^{\star}})z\right)^{2}\right]}\sqrt{\mathbb{E}\left[\zeta(z)^{2}\right]}$
$\leq\sigma\lVert\nabla\hat{f}(x_{0})\rVert\sin\theta(\nabla f^{\star}(x_{0}),\nabla\hat{f}(x_{0}))\sqrt{2}Gd\sigma^{2}$
$=O(\sigma^{3}),$
where we used the inequality (17) in the second inequality, and the assumption that $\lVert\nabla\hat{f}(x_{0})\rVert$ is upper bounded in the last theorem.
We complete the proof by combining (16), (18), and (17).
Appendix C GAAP Implementation Details
In this section, we first introduce an example of the perturbation set $\mathcal{P}er(S_{0})$, then introduce the details of our GAAP algorithm that is used for the perturbation function $\mathcal{P}er(\cdot)$.
C.1 An example
We show an example of $\mathcal{P}er(S_{x_{0}})$ in Table 3. The original input is a question from TriviaQA $S_{x_{0}}=$“Which golfer became only the fifth in history to get both the British and US Open championships in the same year, in 1982? ”.
$S_{x_{0}}$ Which golfer became only the fifth in history to get both the British and US Open championships in the same year, in 1982?
$S_{x_{1}}$ Which golfer became only the fifth in history to get both the British and US Open in the same year, in 1982?
$S_{x_{2}}$ Which driver became only the fifth in history to get both the British and US Open triple_crown in the same year, in 1982?
$S_{x_{3}}$ Which medalist became only the fifth in history to get both the British and US Open high_status in the same year, in 1982?
$S_{x_{4}}$ Which golfer became only the fifth in history to win both the British and US Open championships in the same year, in 1982?
$S_{x_{5}}$ Which golfer became only the fifth in history to win both the British and US Open in the same year, in 1982?
$S_{x_{6}}$ Which driver became only the fifth in history to win both the British and US Open triple_crown in the same year, in 1982?
$S_{x_{7}}$ Which medalist became only the fifth in history to win both the British and US Open title in the same year, in 1982?
$S_{x_{8}}$ Which driver became only the fifth in history to win both the British and US Open championships in the same year, in 1982?
$S_{x_{9}}$ Which driver became only the fifth in history to win both the British and US Open in the same year, in 1982?
$S_{x_{10}}$ Which linksman became only the fifth in history to win both the British and US Open in the same year, in 1982?
Table 3: A question $S_{x_{0}}$ and $10$ perturbed versions of $S_{x_{0}}$ as the output of GAAP.
C.2 Algorithm details
We first introduce some notations. For an input sentence $S_{x_{0}}$, we can denote it in the form of a token (word) series,
$S_{x_{0}}=(t_{1},t_{2},\dots,t_{p}),$
where $t_{1},t_{2},\dots,t_{p}$ denote the sequence of tokens that constitute the input text $S_{x_{0}}$, arranged in their original order. We will then elaborate on each step of GAAP.
C.2.1 Key words selection
For better sampling efficiency, GAAP does not perturb all tokens equally. Instead, we identify the key tokens in $S_{x_{0}}$ first and only perturb these tokens. As a result, we could explore the semantic space more efficiently under the perturbation budget constraint.
We define a function $k(\cdot,\cdot)$ that identifies key tokens within a given text. The function takes two inputs: the first is a text and the second is a ratio indicating the proportion of key tokens to all tokens in this text. $k(S_{x_{0}},r)$ returns a subset of tokens:
$k(S_{x_{0}},r)=\{t_{j_{1}},t_{j_{2}},\dots,t_{j_{q}}\},$
where the indices $\{j_{1},j_{2},\dots,j_{q}\}\subseteq\{1,\dots,p\}$, and the number of selected key tokens are: $q=\text{int}[r\cdot p]$. In GAAP, we use KeyBERT [Grootendorst, 2024] to implement $k(\cdot,\cdot)$.
C.2.2 Initial population generation
Next, we define a replacement function $re(\cdot,\cdot,\cdot)$ that substitutes a specific token in the sequence. The function is defined as follows:
$re(S_{x_{0}},t_{j_{i}},t^{\prime}_{j_{i}})=(t_{1},\dots,t_{j_{i}-1},t^{\prime}_{j_{i}},t_{j_{i}+1},\dots,t_{p}).$ (19)
In GAAP, we choose $t^{\prime}_{j_{i}}$ from a substitution set $\text{SUB}(t_{j_{i}})$. And the substitution set is defined as the union of all possible hypernyms, hyponyms, synonyms, and an empty set denoting word deletion, $\text{SUB}(t_{j_{i}})=\text{hypernyms}(t_{j_{i}})\cup\text{hyponyms}(t_{j_{i}})\cup\text{synonyms}(t_{j_{i}})\cup\{\emptyset\}$.
The initial population of GAAP for the input $S_{x_{0}}$ is defined as the union of the outcomes of all possible single-key token perturbations,
$\text{Pop}_{0}(S_{x_{0}})=\bigcup_{t_{j_{i}}\in k(S_{x_{0}},r)}\bigcup_{t^{\prime}_{j_{i}}\in\text{SUB}(t_{j_{i}})}re(S_{x_{0}},t_{j_{i}},t^{\prime}_{j_{i}}).$
C.2.3 Iterative population update
Then, we introduce an iterative scheme for the perturbed population to evolve. In each iteration, we must follow the next three steps in sequence.
  1. 1.
    Selection: The selection step aims to choose a subset of individuals from the population as parents for subsequent procedures. In GAAP, we design a random selection mechanism where individuals whose semantic meanings are closer to those of the original text $S_{x_{0}}$ will be selected with higher probability.
    In the terminology of genetic algorithms, we define our fitness function as the semantic similarity to $x_{0}$, $a_{\text{Similarity}}(x_{0},\cdot)$. Then, we compute the fitness value $a_{\text{Similarity}}(x_{0},x_{i})$ for $\forall S_{x_{i}}\in\text{Pop}_{t}(S_{x_{0}})$, and use roulette wheel selection [Lipowski and Lipowska, 2012] to choose parents. More specifically, the probability of selecting $S_{x_{i}}$ is:
    $P(S_{x_{i}})=\frac{a_{\text{Similarity}}(x_{0},x_{i})}{\sum_{S_{x_{j}}\in\text{Pop}_{t}(S_{x_{0}})}a_{\text{Similarity}}(x_{0},x_{j})}$
    The set of all selected parent individuals is denoted as $\text{Pa}_{t}(S_{x_{0}})$.
  2. 2.
    Crossover: The crossover step aims to generate new offspring by recombining the segments (i.e., token sub-sequences) of parent individuals.
    The inputs of the crossover operation are two randomly selected parent individuals $S_{x_{A}}$ and $S_{x_{B}}$ from $\text{Pa}_{t}(S_{x_{0}})$. Then, we uniformly randomly sample a crossover point $h$ from $\{1,2,\dots,p-1\}$), where $p$ is the length of the shorter one between $S_{x_{A}}$ and $S_{x_{B}}$. Next, we generate two offspring individuals $S_{x_{A^{\prime}}}$ and $S_{x_{B^{\prime}}}$ as
    $S_{x_{A^{\prime}}}=(t_{1}^{A},\dots,t_{h}^{A},t_{h+1}^{B},\dots,t_{p}^{B}),$
    $S_{x_{B^{\prime}}}=(t_{1}^{B},\dots,t_{h}^{B},t_{h+1}^{A},\dots,t_{p}^{A}).$
    where $t_{i}^{A}$ and $t_{i}^{B}$ represent the $i$-th token of parents $S_{x_{A}}$ and $S_{x_{B}}$, respectively. The set of all generated offspring individuals is denoted as $\text{Off}_{t}(S_{x_{0}})$.
  3. 3.
    Mutation: The mutation operation aims to augment population diversity again by randomly replacing certain tokens in the offspring individuals. For each offspring individual $S_{x_{i}}\in\text{Off}_{t}(S_{x_{0}})$, we randomly select a token $t_{j_{i}}$ for replacement. Then, we uniformly randomly choose a new token $t^{\prime}_{j_{i}}$ from the substitution set $\text{SUB}(t_{j_{i}})$:
    $t^{\prime}_{j_{i}}\in\text{SUB}(t_{j_{i}})=\text{hypernyms}(t_{j_{i}})\cup\text{hyponyms}(t_{j_{i}})\cup\text{synonyms}(t_{j_{i}})\cup\{\emptyset\}.$
    Finally, we generate the mutated individual as
    $S_{x_{i}^{\prime}}=re(S_{x_{i}},t_{j_{i}},t^{\prime}_{j_{i}}),$
    where $re(\cdot,\cdot,\cdot)$ is the replacement function defined in (19). The set of all mutated offspring individuals is denoted as $\text{Mu}_{t}(S_{x_{0}})$.
The new population is formed by combining all the 3 sets above,
$\text{Pop}_{t+1}(S_{x_{0}})=\text{Pa}_{t}(S_{x_{0}})\cup\text{Off}_{t}(S_{x_{0}})\cup\text{Mu}_{t}(S_{x_{0}}).$ (20)
Equation (20) defines an iterative algorithm to update the population for $t=0,1,2,\cdots$. The iterative process terminates when either of the following conditions is met: (1) The number of generations $t$ exceeds a predefined maximum value $Num$: $t\geq Num$, (2) the maximum of fitness values in the population is smaller than a threshold $\delta$: $\max_{i}a_{\text{Similarity}}(x_{0},x_{i})<\delta$.
C.2.4 Perturbation set construction
Unlike standard genetic algorithms that aim to optimize the fitness function to its extreme value, the objective for GAAP is to use populations $\text{Pop}_{t}$ at different generations $t$ to construct a perturbation set $\mathcal{P}er(S_{x_{0}})$. The ideal perturbation set should be diverse and contain texts with varying degrees of similarity to $S_{x_{0}}$. Since earlier generations contain fewer perturbations and later generations involve more perturbations, the generation index $t$ is a natural indicator of similarity. As a result, we generate $\mathcal{P}er(S_{x_{0}})$ by random sampling from populations at different generations.
More specifically, $\mathcal{P}er(S_{x_{0}})$ consists of random samples from populations at generations $t=0,\tau,2\tau,...$, where $\tau\in\mathbb{Z}$ is the sample interval:
$\mathcal{P}er(S_{x_{0}})=\bigcup_{q=0}^{\text{int}\left(\frac{T}{\tau}\right)}\text{Uniform}(\text{Pop}_{q\tau}(S_{x_{0}}))$ (21)
where $\text{Uniform}(\cdot)$ is a function that uniformly randomly selects a subset of individuals from the population. The construction rule (21) ensures that $\mathcal{P}er(S_{x_{0}})$ contains texts with progressively decreasing similarity to $S_{x_{0}}$, as the genetic algorithm evolves toward lower fitness values (i.e., lower similarity).
After the termination condition is met, the perturbation set $\mathcal{P}er(S_{x_{0}})$ is returned as the final output. This set represents a collection of perturbed versions of $S_{x_{0}}$, each with a different degree of perturbation.
Appendix D Settings & Additional Experimental Results
D.1 Experimental setting
Dataset Preprocessing
We sampled 1,000 data points from each dataset for our experiments. No preprocessing was required prior to sampling, except for MMLU, due to its heterogeneous and often non-standard question formats, which are incompatible with our perturbation-based framework. We excluded items that lacked self-contained semantic meaning and thus were unsuitable for perturbation (e.g., “Which of the following statements is true?”), purely mathematical expressions without contextual meaning (e.g., “If $A=(1,2,3,4)$, let $B=\{(1,2),(1,3),(4,2)\}$. Then $B$ is”), or other irregular or ambiguous phrasing. Only questions with clearly worded, self-contained statements were retained.
The following are all the prompts used in our experiments.
ChatGPT-based paraphrasing
Please Provide {number of perturbations} paraphrases for this sentence: {sentence}
Generating Responses
For TriviaQA, SciQ, and NQ:
{question} Answer concisely and return only the name.
For MMLU:
{question + choices} Answer concisely and return only the name.
For GSM8K:
{question} Answer concisely and return only the result itself.
We design our prompts to align closely with the highly concise reference answers in the datasets. The same prompts are used for both our method and all baseline models, across both GPT-3.5-Turbo and LLaMA-3.1-8B-Instruct.
Correctness Evaluation
Are the following two answers to my question Q semantically equivalent?
Q: {question}
A1: {standard answer}
A2: {answer}
Please answer with a single word, either Yes or No.
VU Derivation (used as one of our benchmarks)
After the previous prompt of generating responses, we append the following prompt to elicit verbalized uncertainty:
And use a percentage to tell me your confidence in your answer.
The parameters used in our experiments are listed below.
Perturbation Configuration
As detailed in the main manuscript, we generate nine perturbations per question using ChatGPT-based paraphrasing, resulting in ten variants per question including the original. For GAAP, we set the threshold $\delta=0.7$ and also fix the number of perturbations at nine. This uniformity is crucial for our probabilistic framework, where each question induces a distribution over variants. To ensure that these distributions are comparable and defined on the same scale, we require the same number of perturbations per question across the dataset. When fewer than nine are generated, we randomly duplicate existing perturbations; when more than nine are produced, we randomly sample nine.
Replication and Bootstrapping
Our model incorporates bootstrapping to utilize replicated responses. Unless otherwise specified (notably in the ablation studies analyzing the impact of $S$ and $r$), all reported results are based on experiments with $S=30$ bootstrapping iterations and $r=5$ replications.
Calculation of Mean and Variance
All reported evaluation metrics represent means with associated standard deviations computed via bootstrapping. Using 40 bootstrap samples generated with replacement, we: (i) Calculate the target metric for each sample, (ii) Aggregate results by taking the mean of sample-level metrics as the final estimate, (iii) Compute the standard deviation across the 40 values as a dispersion measure.
LLM Configuration
For ChatGPT-based paraphrasing, we set the temperature to 0.7. For correctness evaluation, a temperature of 0 is used. For LLaMA, due to LLaMA’s lack of automatic response termination and occasional output corruption, this may lead to incomplete or malformed answers. To mitigate this, we adopt a multi-attempt generation protocol with the following cleaning steps to ensure concise and valid outputs: (i) remove the echoed question if present, (ii) delete formatting tokens (e.g., [INST], [/INST], #) and any trailing text, and (iii) retain only the first non-empty line after trimming whitespace. The following is an example.
Question:
What is the capital of France?
Response before cleaning:
What is the capital of France?
Paris [/INST]#
It is a major European city and a global
Response after cleaning:
Paris
D.2 Additional experimental results
Table 4 shows how the experimental results vary with the number of bootstrapping iterations $S$. We observe that increasing $S$ initially enhances the performance of our method, Inv-Entropy. However, beyond a certain threshold, further increases yield diminishing or no significant returns. Therefore, using more than 10 iterations appears to provide limited additional benefit.
Table 5 shows the variation of several proposed UQ measures (Inv-Entropy, NI-Entropy, WD-px-py, MAX-py-x) as the number of replications $r$ increases. Performance improves consistently with larger $r$, as additional replications help better capture aleatoric uncertainty. However, this improvement comes at the expense of increased computational cost.
Table 6 demonstrates the robustness of our framework across different embedding functions. On all three datasets (TriviaQA, SciQ, and MMLU), Inv-Entropy performs strongly with SBERT-small, SBERT-large, and DeBERTa. This consistency allows practitioners to select an encoder based on available computational resources or domain-specific requirements without compromising the quality of uncertainty estimation.
Table 7 and Table 8 present how various uncertainty measures respond to temperature changes on the TriviaQA and SciQ datasets, respectively. Our probabilistic methods consistently outperform baseline models across all temperature settings, highlighting their robustness to decoding variability.
Table 9 presents the results obtained using LLaMA. Inv-Entropy continues to demonstrate strong performance, achieving competitive or state-of-the-art results across most datasets and evaluation metrics.
Table 4: Comparison of Inv-Entropy performance across all the 5 datasets with varying numbers of bootstrapping iterations $S$. We use GPT-3.5-Turbo with ChatGPT-based paraphrasing and DeBERTa-v2-xlarge-MNLI embedding function.
Dataset Metric $S{=}1$ $S{=}5$ $S{=}10$ $S{=}30$ $S{=}50$ $S{=}100$
TriviaQA AUROC 0.743$\pm$0.061 0.781$\pm$0.057 0.785$\pm$0.056 0.788$\pm$0.054 0.780$\pm$0.054 0.788$\pm$0.054
PRR 0.840$\pm$0.054 0.881$\pm$0.043 0.882$\pm$0.043 0.885$\pm$0.044 0.882$\pm$0.044 0.885$\pm$0.044
Brier 0.138$\pm$0.021 0.127$\pm$0.020 0.125$\pm$0.020 0.128$\pm$0.020 0.131$\pm$0.021 0.128$\pm$0.020
SciQ AUROC 0.724$\pm$0.052 0.733$\pm$0.050 0.740$\pm$0.049 0.740$\pm$0.050 0.739$\pm$0.050 0.743$\pm$0.049
PRR 0.821$\pm$0.052 0.840$\pm$0.046 0.844$\pm$0.045 0.843$\pm$0.042 0.842$\pm$0.046 0.845$\pm$0.046
Brier 0.163$\pm$0.016 0.160$\pm$0.015 0.159$\pm$0.015 0.157$\pm$0.018 0.160$\pm$0.015 0.159$\pm$0.015
NQ AUROC 0.659$\pm$0.063 0.686$\pm$0.060 0.699$\pm$0.057 0.703$\pm$0.060 0.702$\pm$0.058 0.705$\pm$0.059
PRR 0.709$\pm$0.068 0.730$\pm$0.071 0.760$\pm$0.065 0.764$\pm$0.064 0.764$\pm$0.063 0.766$\pm$0.064
Brier 0.199$\pm$0.019 0.194$\pm$0.020 0.184$\pm$0.019 0.182$\pm$0.020 0.183$\pm$0.020 0.182$\pm$0.021
MMLU AUROC 0.604$\pm$0.059 0.762$\pm$0.042 0.777$\pm$0.042 0.780$\pm$0.041 0.790$\pm$0.041 0.789$\pm$0.040
PRR 0.743$\pm$0.064 0.863$\pm$0.046 0.897$\pm$0.031 0.898$\pm$0.030 0.905$\pm$0.028 0.902$\pm$0.030
Brier 0.188$\pm$0.021 0.152$\pm$0.017 0.148$\pm$0.017 0.147$\pm$0.017 0.142$\pm$0.017 0.143$\pm$0.017
GSM8K AUROC 0.674$\pm$0.054 0.684$\pm$0.056 0.685$\pm$0.054 0.695$\pm$0.051 0.690$\pm$0.055 0.695$\pm$0.055
PRR 0.487$\pm$0.090 0.530$\pm$0.096 0.527$\pm$0.095 0.521$\pm$0.094 0.520$\pm$0.094 0.527$\pm$0.095
Brier 0.159$\pm$0.021 0.151$\pm$0.022 0.150$\pm$0.021 0.152$\pm$0.020 0.152$\pm$0.020 0.146$\pm$0.021
Table 5: Comparison of UQ methods on the TriviaQA and SciQ datasets with varying numbers of replications $r$. We use GPT-3.5-Turbo with ChatGPT-based paraphrasing and SBERT-large (all-mpnet-base-v2) embedding function.
Dataset TriviaQA SciQ
Metric $r=2$ $r=4$ $r=6$ $r=2$ $r=4$ $r=6$
Inv-Entropy AUROC 0.812$\pm$0.044 0.815$\pm$0.044 0.820$\pm$0.044 0.794$\pm$0.044 0.801$\pm$0.043 0.806$\pm$0.042
PRR 0.916$\pm$0.028 0.915$\pm$0.029 0.915$\pm$0.029 0.888$\pm$0.038 0.888$\pm$0.039 0.892$\pm$0.038
Brier 0.128$\pm$0.020 0.121$\pm$0.020 0.123$\pm$0.020 0.143$\pm$0.019 0.140$\pm$0.019 0.136$\pm$0.019
NI-Entropy AUROC 0.702$\pm$0.072 0.631$\pm$0.081 0.617$\pm$0.083 0.729$\pm$0.056 0.731$\pm$0.058 0.746$\pm$0.055
PRR 0.813$\pm$0.054 0.765$\pm$0.060 0.751$\pm$0.059 0.805$\pm$0.055 0.805$\pm$0.056 0.809$\pm$0.054
Brier 0.133$\pm$0.024 0.152$\pm$0.023 0.150$\pm$0.024 0.147$\pm$0.021 0.146$\pm$0.021 0.144$\pm$0.021
WD-px-py AUROC 0.762$\pm$0.058 0.771$\pm$0.058 0.772$\pm$0.057 0.684$\pm$0.052 0.688$\pm$0.058 0.685$\pm$0.057
PRR 0.866$\pm$0.046 0.875$\pm$0.044 0.876$\pm$0.044 0.825$\pm$0.050 0.818$\pm$0.054 0.819$\pm$0.051
Brier 0.134$\pm$0.021 0.131$\pm$0.021 0.130$\pm$0.021 0.173$\pm$0.017 0.169$\pm$0.020 0.172$\pm$0.019
MAX-py-x AUROC 0.782$\pm$0.047 0.794$\pm$0.045 0.804$\pm$0.044 0.754$\pm$0.049 0.760$\pm$0.051 0.757$\pm$0.049
PRR 0.904$\pm$0.028 0.910$\pm$0.027 0.915$\pm$0.026 0.864$\pm$0.047 0.862$\pm$0.049 0.861$\pm$0.048
Brier 0.134$\pm$0.019 0.130$\pm$0.019 0.124$\pm$0.019 0.154$\pm$0.018 0.155$\pm$0.019 0.154$\pm$0.017
Table 6: Comparison of AUROC, PRR, and Brier scores across the TriviaQA, SciQ, and MMLU datasets using different embedding functions. We use GPT-3.5-Turbo and ChatGPT-based paraphrasing.
Dataset Metric Embedding Inv-Entropy NI-Entropy WD-px-py MAX-py-x
TriviaQA AUROC SBERT-small 0.792$\pm$0.051 0.666$\pm$0.072 0.833$\pm$0.050 0.835$\pm$0.046
SBERT-large 0.772$\pm$0.057 0.617$\pm$0.083 0.804$\pm$0.044 0.816$\pm$0.044
DeBERTa 0.788$\pm$0.054 0.786$\pm$0.057 0.518$\pm$0.060 0.723$\pm$0.054
PRR SBERT-small 0.899$\pm$0.037 0.803$\pm$0.053 0.920$\pm$0.029 0.920$\pm$0.029
SBERT-large 0.876$\pm$0.044 0.751$\pm$0.059 0.915$\pm$0.026 0.915$\pm$0.030
DeBERTa 0.885$\pm$0.044 0.883$\pm$0.043 0.763$\pm$0.051 0.875$\pm$0.038
Brier SBERT-small 0.127$\pm$0.021 0.151$\pm$0.022 0.106$\pm$0.021 0.114$\pm$0.021
SBERT-large 0.130$\pm$0.021 0.150$\pm$0.024 0.124$\pm$0.019 0.124$\pm$0.020
DeBERTa 0.128$\pm$0.020 0.124$\pm$0.020 0.184$\pm$0.019 0.148$\pm$0.019
SciQ AUROC SBERT-small 0.774$\pm$0.049 0.750$\pm$0.054 0.655$\pm$0.058 0.742$\pm$0.050
SBERT-large 0.806$\pm$0.042 0.746$\pm$0.055 0.685$\pm$0.057 0.757$\pm$0.049
DeBERTa 0.740$\pm$0.050 0.681$\pm$0.056 0.303$\pm$0.060 0.674$\pm$0.054
PRR SBERT-small 0.874$\pm$0.044 0.820$\pm$0.059 0.796$\pm$0.054 0.852$\pm$0.049
SBERT-large 0.892$\pm$0.038 0.809$\pm$0.054 0.819$\pm$0.051 0.861$\pm$0.048
DeBERTa 0.843$\pm$0.042 0.781$\pm$0.053 0.587$\pm$0.056 0.821$\pm$0.048
Brier SBERT-small 0.147$\pm$0.021 0.150$\pm$0.021 0.181$\pm$0.019 0.160$\pm$0.018
SBERT-large 0.136$\pm$0.019 0.144$\pm$0.021 0.172$\pm$0.019 0.154$\pm$0.017
DeBERTa 0.157$\pm$0.018 0.164$\pm$0.017 0.212$\pm$0.016 0.177$\pm$0.017
MMLU AUROC SBERT-small 0.689$\pm$0.060 0.576$\pm$0.063 0.723$\pm$0.049 0.704$\pm$0.056
SBERT-large 0.634$\pm$0.058 0.532$\pm$0.057 0.670$\pm$0.055 0.676$\pm$0.057
DeBERTa 0.780$\pm$0.041 0.710$\pm$0.052 0.573$\pm$0.061 0.585$\pm$0.059
PRR SBERT-small 0.812$\pm$0.064 0.719$\pm$0.069 0.869$\pm$0.035 0.832$\pm$0.057
SBERT-large 0.790$\pm$0.059 0.695$\pm$0.068 0.834$\pm$0.049 0.820$\pm$0.056
DeBERTa 0.898$\pm$0.030 0.823$\pm$0.055 0.777$\pm$0.054 0.749$\pm$0.062
Brier SBERT-small 0.170$\pm$0.019 0.192$\pm$0.020 0.163$\pm$0.018 0.165$\pm$0.019
SBERT-large 0.185$\pm$0.020 0.200$\pm$0.020 0.177$\pm$0.019 0.173$\pm$0.020
DeBERTa 0.147$\pm$0.017 0.168$\pm$0.021 0.188$\pm$0.018 0.189$\pm$0.020
[…]
Metric Method Temperature
$\mathbb{T}$=0.3 $\mathbb{T}$=0.7 $\mathbb{T}$=1.0 $\mathbb{T}$=1.4
AUROC($\uparrow$) Semantic Entropy 0.579$\pm$0.044 0.679$\pm$0.052 0.700$\pm$0.062 0.732$\pm$0.051
VU 0.695$\pm$0.060 0.625$\pm$0.058 0.629$\pm$0.063 0.578$\pm$0.067
P(True) 0.604$\pm$0.050 0.592$\pm$0.046 0.571$\pm$0.043 0.583$\pm$0.038
LexSim 0.649$\pm$0.055 0.715$\pm$0.051 0.775$\pm$0.055 0.792$\pm$0.051
DegMat 0.734$\pm$0.056 0.688$\pm$0.067 0.794$\pm$0.049 0.730$\pm$0.059
LUQ 0.637$\pm$0.067 0.712$\pm$0.059 0.748$\pm$0.063 0.817$\pm$0.044
KLE 0.333$\pm$0.054 0.327$\pm$0.059 0.221$\pm$0.056 0.155$\pm$0.042
Inv-Entropy 0.870$\pm$0.044 0.795$\pm$0.051 0.827$\pm$0.043 0.810$\pm$0.042
NI-Entropy 0.762$\pm$0.063 0.666$\pm$0.073 0.634$\pm$0.083 0.673$\pm$0.055
WD-px-py 0.829$\pm$0.044 0.827$\pm$0.051 0.816$\pm$0.054 0.829$\pm$0.039
MAX-py-x 0.854$\pm$0.043 0.832$\pm$0.046 0.856$\pm$0.040 0.846$\pm$0.041
PRR($\uparrow$) Semantic Entropy 0.787$\pm$0.044 0.803$\pm$0.046 0.834$\pm$0.043 0.811$\pm$0.045
VU 0.836$\pm$0.044 0.791$\pm$0.041 0.798$\pm$0.053 0.727$\pm$0.051
P(True) 0.797$\pm$0.042 0.760$\pm$0.044 0.776$\pm$0.042 0.726$\pm$0.046
LexSim 0.810$\pm$0.045 0.824$\pm$0.040 0.877$\pm$0.043 0.854$\pm$0.043
DegMat 0.882$\pm$0.041 0.816$\pm$0.053 0.895$\pm$0.037 0.812$\pm$0.054
LUQ 0.854$\pm$0.043 0.856$\pm$0.042 0.874$\pm$0.048 0.893$\pm$0.039
KLE 0.704$\pm$0.048 0.646$\pm$0.050 0.623$\pm$0.051 0.516$\pm$0.049
Inv-Entropy 0.939$\pm$0.028 0.900$\pm$0.037 0.936$\pm$0.022 0.903$\pm$0.037
NI-Entropy 0.862$\pm$0.043 0.799$\pm$0.054 0.777$\pm$0.060 0.765$\pm$0.055
WD-px-py 0.938$\pm$0.021 0.918$\pm$0.029 0.912$\pm$0.043 0.923$\pm$0.025
MAX-py-x 0.932$\pm$0.029 0.920$\pm$0.030 0.946$\pm$0.021 0.913$\pm$0.039
Brier($\downarrow$) Semantic Entropy 0.166$\pm$0.023 0.150$\pm$0.026 0.141$\pm$0.023 0.156$\pm$0.023
VU 0.160$\pm$0.022 0.178$\pm$0.017 0.175$\pm$0.023 0.203$\pm$0.016
P(True) 0.172$\pm$0.022 0.188$\pm$0.020 0.179$\pm$0.021 0.198$\pm$0.020
LexSim 0.151$\pm$0.024 0.146$\pm$0.021 0.128$\pm$0.024 0.127$\pm$0.021
DegMat 0.140$\pm$0.021 0.149$\pm$0.022 0.115$\pm$0.018 0.145$\pm$0.023
LUQ 0.148$\pm$0.020 0.142$\pm$0.021 0.121$\pm$0.023 0.121$\pm$0.018
KLE 0.188$\pm$0.021 0.199$\pm$0.020 0.192$\pm$0.021 0.218$\pm$0.014
Inv-Entropy 0.085$\pm$0.019 0.127$\pm$0.021 0.117$\pm$0.019 0.132$\pm$0.017
NI-Entropy 0.104$\pm$0.021 0.151$\pm$0.022 0.142$\pm$0.024 0.164$\pm$0.018
WD-px-py 0.115$\pm$0.018 0.102$\pm$0.022 0.117$\pm$0.021 0.121$\pm$0.017
MAX-py-x 0.103$\pm$0.019 0.116$\pm$0.021 0.108$\pm$0.019 0.114$\pm$0.016
[…]
Metric Method Temperature
$\mathbb{T}$=0.3 $\mathbb{T}$=0.7 $\mathbb{T}$=1.0 $\mathbb{T}$=1.4
AUROC($\uparrow$) Semantic Entropy 0.548$\pm$0.040 0.679$\pm$0.045 0.699$\pm$0.046 0.773$\pm$0.043
VU 0.625$\pm$0.061 0.480$\pm$0.060 0.625$\pm$0.050 0.607$\pm$0.065
P(True) 0.566$\pm$0.037 0.522$\pm$0.026 0.531$\pm$0.026 0.545$\pm$0.026
LexSim 0.552$\pm$0.041 0.681$\pm$0.046 0.713$\pm$0.049 0.770$\pm$0.051
DegMat 0.569$\pm$0.065 0.672$\pm$0.059 0.730$\pm$0.055 0.816$\pm$0.040
LUQ 0.565$\pm$0.062 0.740$\pm$0.050 0.650$\pm$0.060 0.772$\pm$0.051
KLE 0.386$\pm$0.054 0.341$\pm$0.056 0.277$\pm$0.059 0.184$\pm$0.051
Inv-Entropy 0.767$\pm$0.047 0.800$\pm$0.044 0.754$\pm$0.046 0.795$\pm$0.046
NI-Entropy 0.682$\pm$0.062 0.755$\pm$0.055 0.596$\pm$0.065 0.649$\pm$0.063
WD-px-py 0.711$\pm$0.053 0.689$\pm$0.056 0.795$\pm$0.047 0.771$\pm$0.045
MAX-py-x 0.766$\pm$0.048 0.760$\pm$0.048 0.796$\pm$0.046 0.793$\pm$0.044
PRR($\uparrow$) Semantic Entropy 0.705$\pm$0.045 0.763$\pm$0.044 0.780$\pm$0.043 0.798$\pm$0.046
VU 0.739$\pm$0.062 0.677$\pm$0.053 0.736$\pm$0.053 0.678$\pm$0.061
P(True) 0.713$\pm$0.044 0.679$\pm$0.050 0.687$\pm$0.045 0.634$\pm$0.045
LexSim 0.701$\pm$0.052 0.770$\pm$0.051 0.794$\pm$0.047 0.796$\pm$0.056
DegMat 0.739$\pm$0.060 0.802$\pm$0.046 0.834$\pm$0.046 0.864$\pm$0.040
LUQ 0.709$\pm$0.058 0.843$\pm$0.042 0.734$\pm$0.061 0.813$\pm$0.061
KLE 0.632$\pm$0.048 0.592$\pm$0.059 0.554$\pm$0.052 0.479$\pm$0.051
Inv-Entropy 0.900$\pm$0.029 0.888$\pm$0.039 0.864$\pm$0.044 0.870$\pm$0.035
NI-Entropy 0.814$\pm$0.047 0.820$\pm$0.055 0.707$\pm$0.059 0.693$\pm$0.064
WD-px-py 0.858$\pm$0.040 0.820$\pm$0.052 0.868$\pm$0.050 0.851$\pm$0.037
MAX-py-x 0.893$\pm$0.029 0.862$\pm$0.048 0.887$\pm$0.037 0.867$\pm$0.034
Brier($\downarrow$) Semantic Entropy 0.203$\pm$0.018 0.173$\pm$0.020 0.171$\pm$0.021 0.159$\pm$0.022
VU 0.191$\pm$0.022 0.196$\pm$0.017 0.202$\pm$0.018 0.212$\pm$0.016
P(True) 0.204$\pm$0.017 0.215$\pm$0.017 0.212$\pm$0.016 0.225$\pm$0.013
LexSim 0.207$\pm$0.020 0.179$\pm$0.020 0.169$\pm$0.022 0.156$\pm$0.019
DegMat 0.198$\pm$0.017 0.164$\pm$0.018 0.149$\pm$0.020 0.135$\pm$0.020
LUQ 0.195$\pm$0.020 0.157$\pm$0.018 0.173$\pm$0.021 0.159$\pm$0.020
KLE 0.216$\pm$0.016 0.218$\pm$0.016 0.219$\pm$0.016 0.233$\pm$0.011
Inv-Entropy 0.151$\pm$0.018 0.139$\pm$0.020 0.153$\pm$0.020 0.146$\pm$0.020
NI-Entropy 0.164$\pm$0.021 0.140$\pm$0.022 0.188$\pm$0.020 0.186$\pm$0.020
WD-px-py 0.167$\pm$0.018 0.175$\pm$0.019 0.140$\pm$0.020 0.156$\pm$0.015
MAX-py-x 0.151$\pm$0.019 0.155$\pm$0.018 0.143$\pm$0.020 0.153$\pm$0.018
Table 9: Comparison of AUROC, PRR, and Brier scores across all 5 datasets. We use LLaMA-3.1-8B-Instruct with ChatGPT-based paraphrasing, and SBERT-large (all-mpnet-base-v2) embedding model. Bold and underline denote the best and second-best performers, respectively.
Metric Method Datasets
TriviaQA SciQ NQ MMLU GSM8K
AUROC($\uparrow$) Semantic Entropy 0.708 $\pm$0.053 0.660 $\pm$0.061 0.612 $\pm$0.068 0.547 $\pm$0.077 0.511 $\pm$0.085
VU 0.613 $\pm$0.052 0.552 $\pm$0.059 0.501 $\pm$0.065 0.437 $\pm$0.073 0.390 $\pm$0.080
P(True) 0.578 $\pm$0.055 0.516 $\pm$0.050 0.458 $\pm$0.044 0.391 $\pm$0.040 0.321 $\pm$0.035
LexSim 0.695 $\pm$0.054 0.658 $\pm$0.058 0.637 $\pm$0.061 0.587 $\pm$0.066 0.534 $\pm$0.070
DegMat 0.726 $\pm$0.049 0.673 $\pm$0.062 0.616 $\pm$0.074 0.558 $\pm$0.088 0.522 $\pm$0.101
LUQ 0.741 $\pm$0.046 0.673 $\pm$0.060 0.597 $\pm$0.073 0.537 $\pm$0.088 0.502 $\pm$0.102
KLE 0.269 $\pm$0.055 0.378 $\pm$0.061 0.489 $\pm$0.066 0.514 $\pm$0.073 0.548 $\pm$0.079
Inv-Entropy 0.796 $\pm$0.042 0.702 $\pm$0.061 0.631 $\pm$0.079 0.601 $\pm$0.099 0.582 $\pm$0.118
NI-Entropy 0.706 $\pm$0.053 0.653 $\pm$0.063 0.586 $\pm$0.072 0.524 $\pm$0.083 0.486 $\pm$0.093
NR-Inv-Entropy 0.771 $\pm$0.045 0.689 $\pm$0.063 0.612 $\pm$0.080 0.552 $\pm$0.099 0.506 $\pm$0.117
WD-px-py 0.683 $\pm$0.047 0.582 $\pm$0.058 0.489 $\pm$0.068 0.390 $\pm$0.080 0.341 $\pm$0.091
MAX-py-x 0.798 $\pm$0.042 0.677 $\pm$0.060 0.557 $\pm$0.077 0.475 $\pm$0.096 0.428 $\pm$0.114
PRR($\uparrow$) Semantic Entropy 0.739 $\pm$0.058 0.667 $\pm$0.055 0.608 $\pm$0.051 0.536 $\pm$0.049 0.460 $\pm$0.046
VU 0.673 $\pm$0.059 0.610 $\pm$0.058 0.562 $\pm$0.056 0.487 $\pm$0.056 0.431 $\pm$0.055
P(True) 0.616 $\pm$0.065 0.602 $\pm$0.056 0.571 $\pm$0.046 0.555 $\pm$0.038 0.551 $\pm$0.029
LexSim 0.696 $\pm$0.059 0.705 $\pm$0.061 0.733 $\pm$0.062 0.731 $\pm$0.065 0.740 $\pm$0.067
DegMat 0.734 $\pm$0.059 0.753 $\pm$0.058 0.787 $\pm$0.056 0.792 $\pm$0.056 0.825 $\pm$0.055
LUQ 0.746 $\pm$0.055 0.767 $\pm$0.057 0.816 $\pm$0.058 0.831 $\pm$0.061 0.820 $\pm$0.063
KLE 0.440 $\pm$0.066 0.556 $\pm$0.071 0.640 $\pm$0.075 0.700 $\pm$0.081 0.718 $\pm$0.086
Inv-Entropy 0.865 $\pm$0.037 0.810 $\pm$0.058 0.839 $\pm$0.078 0.845 $\pm$0.100 0.862 $\pm$0.121
NI-Entropy 0.702 $\pm$0.065 0.761 $\pm$0.057 0.798 $\pm$0.048 0.802 $\pm$0.041 0.811 $\pm$0.033
NR-Inv-Entropy 0.851 $\pm$0.039 0.795 $\pm$0.060 0.808 $\pm$0.080 0.822 $\pm$0.102 0.816 $\pm$0.123
WD-px-py 0.763 $\pm$0.058 0.742 $\pm$0.064 0.764 $\pm$0.069 0.760 $\pm$0.076 0.770 $\pm$0.082
MAX-py-x 0.855 $\pm$0.041 0.770 $\pm$0.056 0.798 $\pm$0.070 0.799 $\pm$0.086 0.803 $\pm$0.101
Brier($\downarrow$) Semantic Entropy 0.176 $\pm$0.017 0.199 $\pm$0.015 0.221 $\pm$0.013 0.246 $\pm$0.011 0.266 $\pm$0.009
VU 0.185 $\pm$0.013 0.282 $\pm$0.014 0.377 $\pm$0.014 0.476 $\pm$0.016 0.558 $\pm$0.017
P(True) 0.230 $\pm$0.012 0.228 $\pm$0.011 0.228 $\pm$0.010 0.231 $\pm$0.009 0.217 $\pm$0.008
LexSim 0.189 $\pm$0.018 0.203 $\pm$0.018 0.212 $\pm$0.017 0.221 $\pm$0.018 0.242 $\pm$0.018
DegMat 0.179 $\pm$0.017 0.189 $\pm$0.019 0.185 $\pm$0.017 0.198 $\pm$0.023 0.219 $\pm$0.025
LUQ 0.173 $\pm$0.016 0.192 $\pm$0.017 0.199 $\pm$0.018 0.211 $\pm$0.019 0.232 $\pm$0.020
KLE 0.243 $\pm$0.009 0.234 $\pm$0.011 0.206 $\pm$0.013 0.208 $\pm$0.015 0.218 $\pm$0.017
Inv-Entropy 0.149 $\pm$0.018 0.171 $\pm$0.019 0.150 $\pm$0.020 0.146 $\pm$0.021 0.149 $\pm$0.022
NI-Entropy 0.181 $\pm$0.020 0.195 $\pm$0.018 0.199 $\pm$0.015 0.215 $\pm$0.014 0.235 $\pm$0.012
NR-Inv-Entropy 0.158 $\pm$0.018 0.177 $\pm$0.018 0.187 $\pm$0.017 0.201 $\pm$0.018 0.222 $\pm$0.018
WD-px-py 0.185 $\pm$0.014 0.203 $\pm$0.013 0.214 $\pm$0.012 0.229 $\pm$0.011 0.245 $\pm$0.010
MAX-py-x 0.152 $\pm$0.019 0.170 $\pm$0.018 0.156 $\pm$0.016 0.157 $\pm$0.016 0.163 $\pm$0.015